OpenStudy (anonymous):

Solve the equation for x. If necessary, round your answer to two decimal places. Do not include "x =" in your answer. x+7=sqrt16+3x+x^2

OpenStudy (anonymous):

square both sides and solve \[(x+7)^2=x^2+3x+16\]

OpenStudy (anonymous):

that's the thing i need help on how to solve it. I got that part down

OpenStudy (anonymous):

-3+7=sqrt16+3*-3+-3^2

OpenStudy (anonymous):

-3+7=4 4=4+3*-3+-3^2 3*-3=-9 4=4+-9+-3^2 -3^2=9 4=4+-9+9 -9+9=0 4=4+0 4=4 so x=-3

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcome :)